In a school of 80 students, 24 are enrolled in algebra, 49 in music, and 11 students are in both algebra and music. How many students are NOT in either algebra or music?
A) 7 B) 15 C) 18 D) 35 E) 62

Found 2 solutions by richwmiller, fasthomework:
Answer by richwmiller(17219)

There are 80 students.
13+38+11=62 enrolled
80-62=18
so that those in both are not counted twice i removed those eleven from both algebra and music and then added them once
18 are not in either algebra or music

Answer by fasthomework(5)

Firstly, we note that 11 students take both music and algebra.
Hence, this means, that out of the 24 students enrolled for Algebra, 11 of the also take music.
24 - 11 = 13.
Therefore, only 13 students only take Algebra.
Also, this means, that out of the 49 students enrolled in music, 11 of the also take algebra.
49-11= 38.
Therefore, only 38 students only take Music.
The total amount of students enrolled in these two classes are:-
Algebra DISABLED_event_only= 13
Music DISABLED_event_only= 38
Algebra and Music = 11
This gives a total of 13+38+11 = 62 students that are enrolled in the two courses.
Since there are 80 students in the school.
We subtract the 62 students in the two courses from the entire student roll.
80 - 62 = 18
Therefore, 18 students are NOT in either algebra or music.